A 100-m long transmission cable is suspended between two towers. If the mass density is 2.01 kg/m and the tension in the cable is 3.00  104 N, what is the speed of transverse waves on the cable?
1
Expert's answer
2020-04-27T09:59:24-0400
"v=\sqrt{\frac{T}{\mu}}"

"v=\sqrt{\frac{30000}{2.01}}=122\frac{m}{s}"